### v1

# MADS: Multi-Agent Dialogue Simulation for Diverse Persuasion Data Generation

Abstract
We propose MADS (Multi-Agent Dialogue Simulation), a scalable framework for generating persuasive multi-turn dialogues via agent self-play. MADS employs three coordinated agents: User Agents simulating diverse persona-driven behaviors, a Dialog Agent executing task-oriented persuasion strategies and an Optimization Agent evaluating and refining dialogue outcomes. We further validate its effectiveness through users’ Chain-of-Attitude (CoA) modeling and dedicated LLMs’ persuasion assessment. This approach enables low-cost generation of training data without human annotation, addressing key industry challenges such as lack of user data, cold-start evaluation difficulties, and prompt inefficiency. Applied to a real-world marketing scenario, MADS significantly improved the persuasion capacity of small LLMs, increasing the organic traffic conversion rate by 22.4% (from 1.83% to 2.24%) , demonstrating clear business value.
MADS: Multi-Agent Dialogue Simulation for
Diverse Persuasion Data Generation
1 Introduction
Persuasive capability is a critical advantage for task-oriented dialogue systems, particularly in domains such as marketing, healthcare, and finance. Enabling conversational agents to influence user decisions—whether for conversion or engagement—has shown practical benefits in customer-facing applications(Jin et al., 2024)(Liu et al., 2025). Recent studies show that LLM-based agents can exhibit superior moral and emotional language performance compared to humans, raising expectations for their deployment in persuasive tasks(Carrasco-Farré, 2024).
Several approaches have been explored to enhance persuasive dialogue generation. PersRFI reduces redundancy and inconsistency in persuasive dialogues through reinforcement learningShi et al. (2021). In multi-agent systems, COOPER coordinates agents in negotiation and persuasion dialogues, yielding more efficient collaborative strategiesCheng et al. (2023), and Cohesive Conversations enhance dialogue realism through multi-agent inconsistency detectionChu et al. (2024). However, many of these approaches depend heavily on real user data or human feedback, limiting their scalability and practicality in cold-start or early-stage deployments. SpeechAgents introduces vocal modality to simulate emotionally expressive personasZhang et al. (2024), but it remains limited in goal-directed, strategic business scenarios such as marketing.
Figure 1: Architecture of the MADS Framework
Other researchers have noted the limited effectiveness of LLM-based agents in cold-start and long-tail business casesBraggaar et al. (2023), and analyzed issues like goal drift in multi-turn LLM dialoguesLaban et al. (2025) . Other studies explore LLMs’ sensitivity to prompt perturbations across diverse task types (Cao et al., 2024)(Zhuo et al., 2024), highlighting the fragility of current prompting strategies in dynamic settings.
In real-world, LLM-based dialogue systems still face three major challenges:
  • Lack of authentic user data: Existing corpora rarely contain multi-turn dialogues with user profile context, making it difficult to support personalized persuasion modeling.
  • Cold-start evaluation difficulty: New systems lack interaction logs, rendering benchmark evaluations unreliable.
  • Low prompt engineering efficiency: Most prompts are mostly manually designed, making them brittle and difficult to generalize across different user roles and contexts.
To address these limitations, we propose MADS (Multi-Agent Dialogue Simulation), a closed-loop framework for simulating structured, multi-strategy persuasive dialogues through agent self-play (Fig.1). This framework supports scalable generation of high-quality training data with minimal human annotation, making it particularly suitable for cold-start scenarios.
2 Related Work
Multi-Agent Dialogue Simulation and User Modeling: Basic user profile attributes (e.g. name, sex, occupation, age, level of education) constructed from standard demographic attributes are widely used in the recommendation system(Zhang et al., 2018)(Zheng et al., 2019). Seminal works show that such tags can be easily sampled from existing statistical datasets or public samples(Ricci et al., 2015)(Li et al., 2016), and can be automatically assigned using simple probabilistic distributions, significantly enhancing the diversity and conversational realism compared to rule-based agents. Generative Agent Simulations (GAS), combines real user interviews with LLMs to simulate complex social structures, albeit at a higher cost(Park et al., 2024). Recent methods like IntellAgent use LLMs to construct user personas for evaluating unseen scenarios(Levi and Kadar, 2025), while genetic algorithms and multi-agent coordination enhance simulation diversity(Cai et al., 2025). Compared to rule-based approaches, LLM-driven simulation offers greater generalizability and efficiencyWang et al. (2024).
LLM-Based Dialogue Evaluation: LLM-based evaluation has emerged as an alternative to manual scoring. Recent work simulates multi-perspective user feedback(e.g., gender, age, political stance) to improve fairness and task alignment(Wan et al., 2024)(Nan et al., 2024). Combined scoring methods can improve diversity and coherence assessment but remain sensitive to prompts and metrics(Sun et al., 2025).For persuasion-specific tasks, MakeMePay [^1] OpenAI (2024). Make-Me-Pay: OpenAI Evals Suite.GitHub Repository. https://github.com/openai/evals/tree/main/evals/elsuite/make_me_pay and PersuasiveToM (Yu et al., 2025) provide standardized evaluation settings for behavioral reasoning and strategic adaptation. However, existing benchmarks still lack coverage of many real-world scenarios and are not easily adaptable to practical constraints(Giudici, 2024).
Personality Tags and Behavioral Diversity: To simulate a wide range of user behaviors, prior studies have proposed personality-informed user modeling based on traits such as MBTI and zodiac signs. Results suggest that integrating personality cues improves both the diversity and personalization of dialogue strategies(Cheng et al., 2025)(Fernau et al., 2022)(Zhao et al., 2025) . Further evidence shows that users with different personality profiles respond differently to persuasive strategies, confirming the theoretical and practical value of diverse persona modeling in persuasive dialogue systems(Paradeda et al., 2020)(Andrews, 2012)(Ait Baha et al., 2023). Although these trait-based personas are simplifications, they serve as proxies to simulating various user behaviors rather than reflecting real psychological profiles.
3 Methodology
Our methodology is grounded in concrete, domain-specific needs. MADS generates training data through multi-agent self-play simulation, which feeds into a platform-based LLM training pipeline, forming a self-optimizing, closed-loop process: Meta Instruction → Simulation → Optimization → Domain-Specific LLMs. This framework enables low-cost, high-fidelity modeling of user interactions tailored to specific domains. As shown in Fig.1, MADS consists of three modules:
User Agents defines structured user profiles within system prompts, to simulate diverse personas with varying personality traits and business contexts.
Dialog Agent engages in multi-turn interactions with selected User Agents, conducting $N$ independent dialogues per profile.
Optimization Agent automates dialogue annotation and prompt refinement via three sub-modules: Summary, Evaluation, and Evolution. The prompts can be found in AppendixA.4.
We conduct a comprehensive evaluation to assess how MADS improves both dialogue diversity and persuasive effectiveness. Our methodology targets two key aspects: 1) the quality of simulated data, with a focus on attitude diversity and richness of persuasive strategies. 2) the downstream impact on model performance via fine-tuning small-scale dialogue models.
3.1 Chain-of-Attitude (CoA)
User attitude change is a critical signal of success in persuasive tasks. Drawing on classic marketing models like AIDA model(Lewis, 1899)(Corporate Finance Institute, 2024) and the Elaboration Likelihood Model (ELM)(Petty and Cacioppo, 1986)(Cialdini, 2001), we define a structured progression of user attitudes and model their transitions as an attitude chain across multi-turn dialogues.
To capture user attitude dynamics during the dialogue process, we constructed a hierarchical state space consisting of 16 attitude states, the full list of states and their descriptions are included in AppendixA.1. This design follows the principle of progressive attitude change, reflecting the psychological transition from initial resistance to eventual acceptance.
Within a fixed set of user profile tags, user attributes are randomly sampled within predefined ranges to generate $N$ User Agent system prompts. Using each persona prompt along with a base Dialog Agent prompt, we simulate multi-turn dialogues, resulting in a collection of multi-turn dialogues: $\mathcal{D}=\{D^{1},D^{2},...,D^{N}\}$.
In the Optimization Agent’s Summary step, an LLM-based classifier $L^{attitude}$ classifies user attitudes at each turn of multi-turn dialogues. The attitude chain for a single dialogue round is represented as: $X^{n}=L^{attitude}(D^{n})\in S$ where $S$ is the set of attitude states, and the set of CoA for simulation of $N$ is: $\mathbf{X}^{N}=\langle X^{1},X^{2},...,X^{N}\rangle$
Next, we employ a first-order Markov model to represent the sequence of user attitude transitions. Two assumptions are made in this context: 1) Changes in user attitude are primarily influenced by the current dialogue content and the immediate interaction experience. 2) Compared to more distant history, the most recent attitude state has the strongest predictive power for the current decision.
The transition probability matrix for CoA is given by:
$T_{ij}=P(X_{t}=s_{j}|X_{t-1}=s_{i})=\frac{N_{ij}}{\sum_{k=1}^{|S|}N_{ik}}$
Where $N_{ij}$ denotes the count of transitions from state $s_{i}$ to $s_{j}$ observed in $\mathbf{X}^{N}$ and $t$ represents the current dialogue turn.
Based on the above modeling, we use average information entropy as a quantitative metric for the diversity of attitude changes in $\mathcal{D}$:
$H(T_{i})=-\sum_{j:T_{ij}>0}T_{ij}$
$H_{avg}(\mathbf{X}^{N})=\frac{1}{|S|}\sum_{i=1}^{|S|}H(T_{i})$
$\log T_{ij}$ is the information entropy for the i-th row. The theoretical range of this metric is $0,\ln(|S|)]$. For $|S|=16$ states, the maximum value is $\ln(16)\approx 2.77$
We also employ Jensen-Shannon (JS) divergence to compare different transition distributions. Suppose the attitude transition distribution of simulated user data generated by MADS is denoted as $D_{mads}$, and for basic user profiles as $D_{base}$. Then,
$JS=\frac{1}{2}D_{KL}(\mathbf{P}_{D_{mads}}||\mathbf{M})+\frac{1}{2}D_{KL}(\mathbf{P}_{D_{base}}||\mathbf{M})$
3.2 Self-Optimizing of Dialog Agent
Algorithm1 presents the workflow of the self-optimizing of Dialog Agent’s strategies by reflection mechanism. Starting from a concise single-turn prompt, the system iteratively generates an improved system prompt and corresponding training dialogues. In each iteration, task-level metrics such as intent compliance rate and CoA quality are dynamically calculated using a modularized LLM evaluation framework (Ramji et al., 2024)(Madaan et al., 2023). High-quality training data can also helps identify and collect long-tail bad cases for further optimization.
In typical recommendation and marketing scenarios, the acceptance rate or the conversion rate is often the key metric(Rashkin et al., 2018). If the User Agent’s acceptance is not clearly defined, simulation tends to default to rejection. We do not aim for unreasonably high acceptance rates, since if the Dialog Agent fully overwhelms the User Agent, the conversation may derail. Therefore, during practical deployment, it is necessary to set an upper bound on acceptance rate, denoted as $\theta$.
Algorithm 1 MADS Self-Optimizing Workflow
[…]
4:for $k=1$ to $K$ do
5:  $D\leftarrow\textsc{GenDialogue}(U,T,P)$
6:  $H\leftarrow\textsc{Summary}(D,T)$ $\triangleright$ Avg entropy
7:  $\tau,DF\leftarrow\textsc{Evaluate}(D,T)$ $\triangleright$ Accept rate
8:  if $\tau\geq\theta\land H\geq H_{\text{Basic}}$ then
9:   return $D,P$
10:  end if
11:  $P\leftarrow\textsc{Evolution}(P,DF)$
12:end for
13:return $D^{*},P^{*}$
3.3 Persuasiveness Improvement
Inspired by the findings of LIMA(Less Is More for AlignmentZhou et al. (2023)), we posit that a small amount of high-quality dialogue data is sufficient for domain-specific scenarios. To validate this, we evaluate how fine-tuning on MADS-simulated data impacts model behavior from two perspectives:
Make Me Pay (MMP) We employ OpenAI’s MMP evaluation, which tests an LLM’s ability to persuade a user to donate money through a multi-turn dialogue. This benchmark assesses aspects such as dialogue guidance, emotional engagement, and the use of persuasive strategies. According to the GPT-4.5 system card222gpt-4.5 System Card. https://openai.com/index/gpt-4-5-system-card/, GPT-4o achieves only a 1% success rate when evaluated on this task, highlights the task’s difficulty.
Persuasion For Good (P4G) We also use the P4G dataset(Wang et al., 2019), which defines ten typical persuasive strategies used in donation scenarios and provides an accompanying classifier. Based on the provided examples, we further refined the strategy descriptions and employed an LLM for strategy classification. When applying a pass@3 criterion (>=2 successes), the classification accuracy of the LLM nearly surpasses that of the original classifier. Leveraging this capability, we extract P4G strategies from the simulated dialogue history $D_{mads}^{N}$, enabling dynamic analysis of the distribution of strategies employed by the Dialog Agent throughout the dialogue generation process.
4 Experiments and Results
4.1 Diversity of CoA and Persuasive Strategy
This experiment aims to quantify the effectiveness of User Agent in MADS for user personality characterization design, and generate high-quality dialogue data with rich attitude changes under the condition of low prompt engineering cost. We designed four groups of User Agent with different portrait labeling system, and Dialog Agent using the original prompt words, to obtain multiple rounds of dialogue collections:
$D_{base}$: user agents with demographic attributes.
$D_{sign}$: adds zodiac signs to $D_{base}$.
$D_{mbti}$: adds MBTI types to $D_{base}$.
$D_{busi}$: adds business attributes to $D_{sign}$.
To visualize differences in attitude dynamics between groups, we plot the normalized transition probability matrices for each group(Fig.2).
Figure 2: Visualization of attitude transition matrices across dialogue collections of different user agent persona groups, (a) $D_{base}$, (b) $D_{sign}$, (c) $D_{mbti}$, (d) $D_{busi}$
As shown in Table 1 , incorporating personality traits such as Zodiac Signs and MBTI into the user profile prompts significantly improves the diversity of simulated user behaviors, which suggesting richer and more nuanced behavioral trajectories across attitude states. From the perspective of Jensen-Shannon divergence, the $D_{busi}$ group shows the highest JS score (0.3442), indicating that the business-oriented profile prompts induce the most distinct CoA distribution compared to the baseline.
Group $H_{\text{avg}}$ $H_{\text{norm}}$ JSJS
$D_{base}$ 1.2982 0.4687 -
$D_{sign}$ 1.7577 0.6345 (↑ 35.37%) 0.2709
$D_{mbti}$ 1.8142 0.6549 (↑ 39.73%) 0.2759
$D_{busi}$ 1.6477 0.5948 (↑ 26.90%) 0.3442
Table 1: Average attitude entropy ($H_{\text{avg}}$) and normalized entropy ($H_{\text{norm}}$) for different user persona groups
We next examine how the diversity of user persona impacts the persuasive strategies employed by the Dialog Agent. During the testing process, we recorded the average number of persuasive strategies used in each simulated conversation. Table 2 below summarizes the overall metrics and presents a comparative analysis between groups.
Metric $D_{base}$ $D_{sign}$ $D_{mbti}$ $D_{busi}$
$C_{str}$ 1.8 2.5 2.7 2.3
Table 2: Average count of distinct persuasive strategies employed by the Dialog Agent
Our analysis shows that increasing the diversity of user personality profiles significantly enriches the persuasive strategies adopted by the Dialog Agent. We also observed that incorporating Zodiac Sign and MBTI attributes leads to a notable increase in the average number of strategies used per dialogue. Moreover, the distribution of strategy types becomes more balanced compared to the baseline, indicating improved coverage and reduced redundancy.
4.2 Data Augmentation via MADS
To verify the effectiveness of simulated data in enhancing model persuasiveness, we constructed synthetic dialogue data under an insurance scenario and used it to finetune small models. The fine-tuned models were then evaluated using the Claude-3.5 and MMP to assess changes in persuasive performance post-finetuning.
Model Donation (%) Withdraw (%)
Original MADS Original MADS
GPT-4o 36 - 34 -
Mistral-7B 14 30 (↑) 76 56 (↓)
ERNIE-Lite 18 30 (↑) 70 58 (↓)
Qwen2.5-7B 40 46 (↑) 44 30 (↓)
Table 3: Performance Comparison on MMP
We evaluated donation success and user withdrawal rates, as shown in Table 3. Fine-tuning with simulation data from insurance scenarios significantly improved performance on the MMP, highlighting the value of offline-generated synthetic training data. The fine-tuned ERNIE-Lite model increased the donation success rate from 18% to 30% and reduced the user withdrawal rate from 70% to 58%. Similar improvements were observed for the open source Qwen2.5 and Mistral models, demonstrating the broad applicability of simulation-based training across model architectures.
4.3 Performance in Simulated Scenarios
MADS based on hierarchical user information and task descriptions. For cases where the task remains incomplete, a reflection mechanism is introduced to analyze and optimize the system prompt. After 2–3 iterations, we observe substantial improvements in task completion rates.
In the marketing scenarios (Table 4), we generated simulated dialogues and observed the following: during the first iteration, not all bad cases are covered, resulting in a limited or even negligible improvement of the Self-Optimizing. After the second iteration, the optimized system prompt demonstrated notable improvements over the initial Meta Instruction (Original Input). Thus, Table 4 reports only the success rates of the initial prompt and the second iteration:
Scenario
Meta Instruction
Success Rate
Prompt@K=2
Success Rate
Automotive 32.5% 45%
Insurance 12.5% 25%
Finance 17.5% 22.5%
Table 4: Prompt Success Rate Comparison across Marketing Scenarios
4.4 Performance in the Real-World Scenario
We trained an end-to-end Audio LLM specifically for the insurance scenario. Compared to the conventional Agent+TTS pipeline, the MADS-trained Audio-LLM-16b model achieved consistent improvements across multiple operational metrics, including organic traffic conversion rate, user engagement, and dialogue length. The results (Table 6.) suggest that applying the MADS framework to end-to-end audio model training enhances overall task effectiveness in the real-world scenario.
Metric Baseline MADS
Conversion Rate (%) 1.83 2.24 ($\uparrow$22.4%)
User Intention Rate (%) 4.53 5.82 ($\uparrow$28.5%)
Avg. Dialogue Turns 1.54 2.06 ($\uparrow$33.8%)
Table 5: Real-world Performance: Agent+TTS (Baseline) vs MADS-Audio-LLM-16b based on a 30,000-sample dataset
Setting Std Dev($\sigma$) CV
Baseline 0.1491 1.3416
Zodiac 0.1012 0.9110
MBTI 0.1122 1.0097
Table 6: Comparison of Variability ($\sigma$ and CV) Across Baseline, Zodiac, and MBTI Settings
Conclusion
Effective persona modeling is essential for realistic user simulation with LLMs. The MADS framework, grounded in Chain-of-Attitude (CoA) modeling, demonstrates that symbolic traits such as zodiac signs and MBTI types can significantly enhance the diversity of simulated dialogue data. In persuasive dialogue tasks, even when prompts remain unchanged., varied persona inputs lead to more diverse and structurally rich persuasion strategies.
To improve performance for specific user segments, MADS employs a self-optimizing mechanism that automatically generates personalized prompts, resulting in higher persuasion success rates. By combining large-scale heterogeneous multi-turn dialogues from simulated user agents with automatic strategy annotation and prompt customization, we fine-tuned and evaluated multiple small-parameter LLMs. Experimental results confirm the effectiveness of MADS-generated data in boosting persuasive performance.
Finally, MADS has been deployed in real-world business settings to customize specific-domain models. Compared to traditional agent-based solutions, our approach achieved over a 28% improvement in user intent rate.
Limitations
1. Single-Dimension Evaluation Limitation
While this work focuses on enhancing persuasive effectiveness in task-oriented dialogue, persuasion represents only one dimension of dialogue quality. A comprehensive evaluation should also account for other aspects, such as factual accuracy, emotional appropriateness, personalization, coherence, and ethical alignment. The absence of multi-dimensional assessment in our current study may limit the completeness of the conclusions drawn, and future work could benefit from incorporating a broader set of evaluation criteria to more holistically measure dialogue system performance.
2. Representational Bias from Fixed Attitude Taxonomy
The attitude chains in this study are constructed based on a manually defined taxonomy with fixed categories. While this structured design facilitates systematic modeling and analysis, it may introduce representational bias by constraining user behavior within a predefined and potentially oversimplified space. Future work could explore data-driven or adaptive attitude representations to capture a broader spectrum of user intent and variability.
3. Prompt Optimization and User-Type Stratification
In the current MADS framework, prompt evolution is performed by simulating multiple user profiles within a given scenario, and refining the prompt based on aggregated feedback across these diverse users. However, this design assumes that a single prompt strategy can effectively accommodate a wide spectrum of personas. In practice, users with different psychological traits (e.g., an INTP Scorpio vs. an ESTJ Leo) may respond positively to entirely different persuasive strategies, making a one-size-fits-all prompt suboptimal or even misleading.
This limitation suggests a more layered approach in the future: first, broadly simulate dialogues across a diverse user population to observe emerging behavioral patterns; then, cluster users based on response or strategy preference; finally, perform stratified prompt evolution within each cluster. Such hierarchical optimization would better capture intra-group coherence and inter-group diversity, leading to more robust and transferable prompt strategies across user types.
4. Discrepancy Between Automatic Metrics and Human Judgment
Current evaluation primarily relies on automatic metrics, such as diversity scores, entropy, and similarity-based clustering. While these are useful for scalable benchmarking, they may not align with human judgment of dialogue quality and persuasive success. The absence of expert or crowd-sourced human evaluation leaves a gap in validating the practical effectiveness of the proposed system.
5. Dependency and Adaptation Challenges with Real-World User Profiling Systems
The current MADS framework operates based on synthetic or predefined user profiles to drive prompt adaptation and dialogue simulation. In practical applications, however, user tags and profiles are often generated by mature recommender or CRM systems using heterogeneous taxonomies (e.g., interest labels, behavioral scores, persona segments). MADS does not need to reinvent these systems, but its performance and applicability are highly dependent on its ability to interface with them. In particular, for users not yet covered by existing tags, it remains unclear whether the upstream pipeline (outside of MADS) can provide sufficient classification granularity in real time. This raises the question of how well MADS can generalize or adapt without reliable profile grounding.
Appendix A Appendix
A.1 MADS Psychological State Space
Table 7 illustrates the hierarchical structure of user attitudes in MADS, aligned with the AIDA model and classical theories of persuasive communication. We define four coarse-grained psychological states—Negative, Neutral, Positive, and Acceptance—corresponding to the Attention, Interest, Desire, and Action stages of AIDA. Each state comprises representative user attitudes commonly observed in persuasive dialogues.
AIDA Stage MADS Psychological State MADS Typical Attitude
Attention Negative Refusal, Resistance, Disinterest, Doubt
Interest Neutral Indifference, Cautious, Hesitant, Weighing Options
Desire Positive Interested, Attention, Consideration, Seeking Value
Action Acceptance Active, Cooperative, Satisfied, Acceptance
Table 7: Mapping of MADS psychological state categories to example user attitudes, inspired by marketing models (AIDA) and persuasion literature
A.2 Self-Optimizing Evaluation Metric
Drawing on design principles from works such as MMRole(Dai et al., 2024), we define a set of evaluation metrics to assess the quality of simulated dialogues (Table 8). These metrics are scored by a LLM-based evaluator. Each score ranges from 0 to 3, except for Task Success, which is binary. Based on the requirements of specific business scenarios, we filter low-quality dialogues using the mean and quantile of evaluation scores.
Role Metric Scale Description
User Agent Authenticity 0-3 How realistic and human-like the user agent’s behavior is
Dialog Agent Relevance 0-3 How well responses match user input and context
Consistency 0-3 How consistently the agent maintains persona and factual accuracy
Efficiency 0-3 How concisely the agent progresses toward task goals
Human-likeness 0-3 How natural and human-like the agent’s conversation style is
Task Success True/False How effectively the agent completes the main tasks
Table 8: LLM-based Evaluation Metrics
A.3 P4G Results Distribution
In the section Diversity of CoA and Persuasive Strategy, we further visualize the distribution of persuasive strategy types across different user persona groups using a stacked bar chart, as shown in Fig.3.
Figure 3: P4G Results Distribution of persuasive strategy types used across different user persona groups
A.4 Prompts of MADS
To facilitate reproducibility and clarify the construction of simulation inputs, we provide representative system prompts. These prompts serve as the initial configurations for the User Agent, Dialog Agent, and Classifiers in our MADS framework.
Prompt Template – User Agent <user_profile> <basic_info> {user_profile} {business_attributes} </basic_info> <personality> Based on the commonly accepted and general {personality_traits} personality traits, please characterize this person’s personality and reflect it in the dialogue. </personality> </user_profile> <task> Based on the above personal profile, please role-play as this person. Note: 1.You are a real person answering the phone. Keep your responses as brief as possible. 2.Make your replies as conversational and colloquial as possible. 3.Only output the dialogue. Do not output any thoughts, inner monologue, or any unnecessary content. </task> Start the task!
Prompt Template – Dialog Agent <reference_information> Product metadata {product_metadata} </reference_information> <task> {prompt_k} </task> Start the task!
Summary Prompt – User CoA Classifier <core_task> Analyze the evolution of user attitudes towards specific events/topics in the provided conversation history. Must analyze attitudes for EVERY round of dialogue, without omitting any round, and each round can only have ONE attitude. Also determine if the user ultimately shows intent to accept the sales pitch (such as leaving contact information, inquiring about details, scheduling viewing appointments, or expressing purchase interest. </core_task> <predefined_attitude_states> {attitude_description} </predefined_attitude_states> <analysis_requirements> Round-by-Round Analysis: Must analyze user attitude in every round of dialogue Single Attitude: Only one primary attitude state can be selected per round State Restriction: Strictly select from predefined set, no other vocabulary allowed Evidence Support: Each attitude judgment must have specific textual evidence Sales Outcome Assessment: Clearly determine if user accepts the sales pitch (Acceptance criteria: clearly scheduling time and place, agreeing to purchase, or making other specific commitment behaviors) </analysis_requirements> <input_source> {dialog_history} </input_source>
Summary Prompt – P4G Classifier <task_description> Extract the persuasion strategies used by the assistant in the conversation based on the Persuasion Strategy Definition Table. Output in JSON format a list of all strategy codes used throughout the conversation history, along with the reasoning for each strategy selection and where they appear in the original text. Note: Do not force-map or extract strategies. Only extract a strategy when its usage fully satisfies the Strategy Description. </task_description> <strategy_definition_table> {strategy_description} </strategy_definition_table> <output_format> { "total_strategies": [{"List of strategy codes"}], "strategy_details": [ { "strategy_id": {"Strategy code"}, "strategy_name": {"Strategy name"}, "reason": {"Explanation for choosing this strategy"}, "occurrences": [ { "turn": {"Conversation turn number"}, "original_text": {"Original text excerpt"}, "explanation": {"How this excerpt demonstrates the strategy"} } ] } ] } </output_format> <input_source> {dialog_history} </input_source> <additional_instructions> Only include strategies when they clearly match the strategy description criteria. Provide detailed reasoning and specific text examples for each identified strategy. </additional_instructions>
Prompt Template – Optimization Agent <task_description> Based on the conversation history, conduct a deep reflection and optimize the Assistant System Prompt. </task_description> <analysis_framework> 1. Task Achievement Assessment - [Success Cases] If expected goals were met, identify key success factors and integrate them into business SOP - [Failure Cases] If goals were not achieved, analyze root causes and develop specific improvement measures 2. Strategy Effect Review - Effective Strategies: Which conversation techniques, pacing, and methods produced positive results? - Areas for Improvement: What aspects were lacking and how did they manifest? - Customer Response: At which points did customers show interest or resistance? 3. Optimization Action Plan - Propose executable improvement solutions for identified issues - Consider differentiated handling strategies for various customer types </analysis_framework> <output_requirements> Based on the above analysis, output a complete Updated Business SOP in the following format: Business SOP Step 1: {Phase Name} Objective: {Specific goal description} Strategic Points: • {Specific executable strategy 1} • {Specific executable strategy 2} • … Key Notes: {Critical reminders} Step 2: {Phase Name} [Continue with similar format…] </output_requirements> <constraints> - All strategies must be based on background knowledge to ensure information accuracy - SOP content should be actionable, avoiding vague descriptions - Prioritize retention of verified effective strategies </constraints> <reference_information> Background {background} Assistant System Prompt {prompt_k} Chat History {dialog_history} </reference_information> <final_instruction> Get Started with the analysis and optimization! </final_instruction>